The general solution of y000 −4 y00 + y0 + 6 y = 0 is: (Hint: r = 2 is a root of the characteristic equation)
Answer #1

Solution:

\small {y}'''-4{y}''+{y}'+6y=0

The characteristic equation is

\small r^{^{3}}-4r^{2}+r+6=0

Using synthetic division we have,

\small 2 \small 1 \small -4 \small 1 \small 6
\small 2 \small -4 \small -6
\small 1 \small -2 \small -3 \small 0

\small \therefore \left ( r-2 \right )\left ( r^{2}-2r-3 \right )=0

\small \therefore \left ( r-2 \right )\left ( r^{2}-3r+r-3 \right )=0

\small \therefore \left ( r-2 \right )\left [ r\left ( r-3 \right )+1\left ( r-3 \right ) \right ]=0

\small \therefore \left ( r-2 \right )\left ( r-3 \right )\left ( r+1 \right )=0

\small \therefore r=-1,2,3

\small \therefore The general solution of the given differential equation is

\small y=c_{1}e^{-x}+c_{2}e^{2x}+c_{3}e^{3x}